A body of mass 3M at rest explodes into two pieces of mass M and 2M. What is the ratio of the kinetic energy of M to that of 2M? (Ans:2)


1
Expert's answer
2022-04-27T13:34:37-0400

"k=\frac{E_2}{E_1}=\frac{\frac{2Mv^2}2}{\frac{Mv^2}2}=2."